please help me with this problem.
Solve using the five-step problem solving process. Show all steps necessary to arrive to your answer.
A 4 ft. by 7 ft. picnic table is placed on an outdoor patio measuring x ft. by (x+3) ft. Find a polynomial for the remaining open space on the patio (in sq. in.)

Answer by mananth(16949)

the length of patio = (x+3) feet = 12(x+3) inches
the width is x feet so width = 12x inches.
..
width of table = 4 ft. = 48 inches
length = 7 ft. = 72 inches
.
Area of patio - area of table = remaining area
..
12x*12(x+3)- 72 *48 = remaining area
144x^2 +432x-3456 remaining area